OpenStudy (anonymous):

How much heat energy is required to raise the temperature of 0.367 kg of copper from 23.0 C to 60.0 C ? The specific heat of copper is 0.0920 cal/(g x C) If 125 cal of heat is applied to a 60.0 g - piece of copper at 23.0 C, what will the final temperature be? The specific heat of copper is 0.0920 cal /(g x C)

OpenStudy (anonymous):

To both of these you need to know the formula \[q=ms \Delta T\] Where: q= energy m=mass s= specific heat of substance \[\Delta T \] =change in temperature. Notice that the specific heat of Cu is given in cal/g x C. This means our answer will be in calories and that we need to convert 0.367 kg to grams. 0.367 kg (1000g/1kg) = 367 g Now we just plug and chug. q=367 x 0.0920 x (60.0-23.0) q= 1249.27 cal since there are 3 sig figs the answer should be 1250 cal B. is a similar concept using the same equation but looking for a different variable (Delta T) q= 125 m=60.0 s=0.0920 Delta T =? Solve for Delta T 125=60.0 x0.0920 x (delta T) (delta T) = 22.6449 or 22.6 (sig figs) We know know that the temperature rose 22.6 degrees from the initial temp (23) Adding these will give you the final temp. 22.6 + 23= 45.6 C